# Online Intrinsic Rewards for Decision Making Agents from Large Language Model Feedback

### v3

@@ 1 Introduction
Reward functions are central to reinforcement learning (RL), and are often assumed to be given as part of the problem definition (Sutton & Barto, 2018). These functions are written to describe the task at hand, and often involve tradeoffs between ease of task definition and ease of policy optimization. For example, assigning a reward of $+1$ for solving the task and 00 otherwise is simple to define and accurately reflects the task goal, but is difficult to optimize due to providing zero gradients almost everywhere.
@@
In this work, we present ONI, a distributed online intrinsic reward and agent learning system — Figure 1 overviews the high-level components. ONI removes the dependency on external datasets beyond the agent’s own experience and enables large-scale RL training with ease. ONI assumes access to captions of observations, similar to previous work Klissarov et al. (2023); Chu et al. (2023). The captions of collected observations are annotated online by an asynchronous LLM server, and both the policy and intrinsic reward model are simultaneously updated using the LLM’s feedback. Such a learning framework allows us to easily instantiate different algorithmic choices for synthesizing LLM feedback. Specifically, we explore three methods: the first one is retrieval-based and simply hashes the annotations; the second builds a binary classification model to distill the sentiment labels returned by the LLM; and the third sends pairs of captions to the LLM server for preference labeling and learns a ranking model, similar to Motif. We demonstrate that ONI is able to match Motif’s performance across a range of challenging, sparse rewards from the NetHack Learning Environment (NLE) (Küttler et al., 2020), solely using the agent’s gathered experience in a single, unified process. We also open-source our code to facilitate future work on LLM-based intrinsic motivation in settings where large offline datasets are not available.
@@ 2 Background
A key research question is how to define or learn the intrinsic reward $r^{\mathrm{int}}$. A first option is to manually define $r^{\mathrm{int}}$ based on task-specific goals, for example a measure of the distance between the agent’s current state and the goal state. However, handcrafting the intrinsic reward function can require significant domain knowledge and must be redone for each new task. A second option is to define $r^{\mathrm{int}}$ to measure some notion of observation novelty, which encourages the agent to systematically explore the environment. This can work well in smaller environments, but fails in ones that cannot be exhaustively explored in a tractable amount of time. A third class of methods, which we focus on in this work, leverage LLMs to automatically synthesize $r^{\mathrm{int}}$ to reflect prior knowledge about the task. We discuss all three classes of methods in more detail in Section 4.
@@ 3.1 System design: distributed PPO with LLM annotations
Our core system, illustrated in Figure 2, is built on top of the Sample Factory library v1.0 (Petrenko et al., 2020) and their asynchronous variant of proximal policy optimization (Schulman et al., 2017), referred to as APPO. APPO operates on a single machine and concurrently runs many environment instances while asynchronously updating policy and value estimates with the usual PPO rules, and adequately handles policy staleness and data transfers between these components. Concretely for NetHack, by running 480 environment instances APPO collects approximately 32k environment interactions per second on a Tesla A100-80GB GPU with 48 CPUs.
@@
For online reward learning via LLM annotations in this system111 The prior work by Klissarov et al. (2023) in the offline setting was able to connect their intrinsic reward function into Sample Factory’s APPO implementation with a few lines of code, as it just needed to load the PyTorch module of the learned reward function., we added 1) an LLM server hosted on a separate node, 2) an asynchronously running process that passes observation captions to the LLM server via HTTP request, 3) a hash table that stores the captions and corresponding LLM annotations, and 4) learning code that dynamically updates a reward model that learns on these annotations. Without being asynchronous, these components have the potential to block the concurrent execution and can reduce the throughput of the overall system, because calling into an LLM and updating the reward model are time-consuming. We added them in a way that retains most of the throughput (!), approximately 80-95% of the original: the average throughput is 30k environment interactions per second if we do not train any additional reward model to distill the LLM annotations, and 26k if a classification-based reward model is learned at the same time (see Section 3.2 for the description of those approaches). The throughput of LLM annotation depends on the actual LLM and prompt we are using. For the reader’s reference, when hosting a LLaMA-3.1-8B server on a Tesla A100-80GB GPU using the prompt displayed in Appendix B, ONI annotates approximately 810k NetHack captions over the whole training process (2B environment steps, 19-20 hours). The rest of this section overviews the relevant components and how we have modified them.
@@
Our New Asynchronous LLM Worker and Remote LLM Server This worker has input and output queues that communicate back with the learner process, which we use in a way that does not block the main execution of the system. The LLM worker awaits new observations to label from the input queue, formats them into prompts, calls into an LLM, and returns the annotations in the output queue. Additionally, the prompt templates and other LLM options are configured by the LLM worker. We use an annotation and message format that support all the intrinsic reward labels that we consider in Section 3.2. As Sample Factory already utilizes most of the free CPU and GPU capacity of the system, we opted to call into the LLM via an HTTP/REST interface rather than loading it in this process directly. The communication here adds minimal overhead, and also avoids the need to coordinate the shared memory of multiple GPUs between the main APPO code and LLM. Alongside every APPO run, we allocate an unloaded machine with a new instance of vLLM (Kwon et al., 2023b) to exclusively serve the queries from the RL run. Our system also allows using the same node for both processes if computational resources allow, which further reduces communication costs.
@@
To prevent overfitting the limited amount of annotations in the early stage of training, the training thread only continuously updates the reward model after we receive $25$k annotations. Before that, the feedback processing thread runs a few updates of the reward model every time we receive new annotations. Now that we have a flexible way of annotating messages and fitting a reward model on top of them, we turn to defining the types of intrinsic reward functions.
@@ 3.2 Intrinsic Reward Functions
If $c_{t}$ is unlabeled, it is placed into a last-in-first-out (LIFO) queue $Q$ managed by the LLM annotation process, and then sent to the LLM server. The LLM continuously processes elements $c_{i}$ from $Q$ and returns their labels $y_{i}$ to the data processing thread of the learner worker, where the pairs $(c_{i},y_{i})$ are added into $\mathcal{H}$. As described in Section 3.1, the training thread and the feedback processing thread run asynchronously. Therefore, editing the hash table does not slow down policy training. This retrieval-based approach does not generalize to observations with unlabeled captions. However, the resulting intrinsic reward is simple and hyperparameter-free, and may work well when the set of captions belongs to a relatively small set.
@@
where $\mathbb{I}$ is the indicator function. We study the impact of $\eta$ in Section 5.3. Unlike the previous approach, this method has the potential to generalize to observations whose captions are similar, but not identical, to the captions labeled by the LLM. However, like the previous approach, it will assign a same reward to observations which are slightly positive (such as finding a few gold pieces in NetHack) and very positive (finding hundreds of gold pieces or a rare artifact).
@@
where we use average log-likelihood when $y=\varnothing$ [^3] In the equivalent cross entropy minimization formulation, this amounts to using a uniform target $[\frac{1}{2},\frac{1}{2}]$ when $y=\varnothing$. and use the Bradley-Terry model Bradley & Terry (1952) $P_{\phi}(o_{1}\succ o_{2})=\exp\big{(}r^{\mathrm{int}}_{\phi}(o_{1})\big{)}/\big{[}\exp\big{(}{r^{\mathrm{int}}_{\phi}(o_{1})}\big{)}+\exp\big{(}r^{\mathrm{int}}_{\phi}(o_{2})\big{)}\big{]}$. Motif computes the mean $\mu_{\mathcal{D}}$, standard deviation $\sigma_{\mathcal{D}}$, and a fixed quantile $\nu_{\mathcal{D}}$ of $r^{\mathrm{int}}_{\phi}$ over the offline dataset of annotations $\mathcal{D}$. During RL training, it normalizes and thresholds $r^{\mathrm{int}}_{\phi}$ to give the reward
@@ 4 Related Work
Goal-conditioned Reward Design A different approach to reward function design is to define rewards as the distance between the agent’s current state and the goal. For example, one line of work learns a state embedding in a self-supervised manner which converts geodesic distances in the original space to Euclidean distances in feature space (Wu et al., 2019; Wang et al., 2021; Gomez et al., 2024). Another line of work of (Fan et al., 2022; Rocamonde et al., 2023; Adeniji et al., 2023; Kim et al., 2024) leverages pretrained image and text encoders, and defines rewards to be some measure of similarity between embeddings of visual observations and embeddings of textual task descriptions. Using a question generation and answering system, Carta et al. (2022) extract auxiliary objectives from the goal description and construct intrinsic rewards. An interesting combination of goal-conditioned and LLM-aided reward design is the ELLM approach introduced in Du et al. (2023b), which generates candidate goals and uses the distance in LLM embedding space to define the reward. This approach shares the limitations of the works of Kwon et al. (2023a); Chu et al. (2023) discussed in the previous section, in that it requires an LLM call for each agent observation, which becomes computationally infeasible in high-throughput settings involving billions of observations. We discuss more works that utilize LLM for RL broadly in Appendix D.
@@ 5 Experiments
Methods and Hyperparameters We instantiate ONI with the three reward functions described in Section 3.2. We name the three approaches ONI-retrieval, ONI-classification and ONI-ranking, respectively. For policy learning, we use the Chaotic Dwarven GPT5 (CDGPT5) architecture (Myffili, 2021) used in prior work (Piterbarg et al., 2023; Kurenkov et al., 2023; Klissarov et al., 2023). Architecture details can be found in Section A.1. All the methods are trained with two billion $(2\times 10^{9})$ environment steps. We train ONI-classification with with classification threshold $\eta=0.5$, where the intrinsic reward coefficient is $\beta=0.1$ for the Score task and $\beta=0.4$ for all the other sparse-reward tasks and the reward-free agent. Similarly, ONI-retrieval uses $\beta=0.1$ for Score and $\beta=0.5$ for the others. ONI-ranking is trained with $\beta=0.05$ and $\nu_{\mathcal{N}}=0$. Full details of the training process can be found in Section A.2.
LLMs We use the LLaMA-3 herd of models (Dubey et al., 2024) as our LLMs. All prompts are listed in Appendix B. We initially reran the Motif baseline using the official code [^5] https://github.com/facebookresearch/motif and compared the performance of LLaMA-3.1-70B-Instruct and LLaMA-3.1-8B-Instruct (see Section C.1). We did not observe a significant difference in their performance on the Oracle or Score tasks. Therefore, we use LLaMA-3.1-8B-Instruct in our subsequent experiments to reduce computation.
Baselines We compare to 3 baselines: i) agents trained with extrinsic reward alone, ii) Motif (Klissarov et al., 2023), and iii) a variant of the ELLM algorithm Du et al. (2023a). It is not feasible to directly apply ELLM to our setting, since it requires an LLM call for each observation, where the total number of calls scales to the billions in our case. Therefore, we designed a more scalable variant which i) replaces the LLM embedding with a lightweight bag-of-words embedding  [^6] We also tried using Sentence Transformers (SBERT) for generating sentence embeddings, which led to an approximately 20x throughput drop: 28k vs 1.4k FPS on a Tesla V100-32GB GPU, when the sentence embeddings can be cached. If we recompute the sentence embedding in the forward pass every time, the throughput drop is approximately 100x. using FastText (Bojanowski et al., 2016), and ii) includes an episodic term in the intrinsic reward, as with Motif and ONI (see Section A.2). We call it ELLM-BoW and include more details in Appendix A.3. We note that Klissarov et al. (2023) found that other exploration methods based on novelty bonuses such as RND (Burda et al., 2019), NovelD (Zhang et al., 2021) and E3B (Henaff et al., 2022a) did not improve over the extrinsic reward baseline on these tasks, hence we do not include them.
@@ 5.1 Main Results
Task Performance We report the average performance and $95\%$ confidence intervals computed via standard errors over $5$ seeds in Figure 3(a). The extrinsic reward agent performs reasonably well on the dense Score task, but completely fails on the others due to reward sparsity. We find that, despite not requiring any external data, our ONI-classification agent is able to match Motif on all tasks except Oracle, where it still performs well. We note that Motif requires pre-collecting data from the Score task even for sparse reward tasks—this assumes access to an additional dense reward function, which is an often unrealistic assumption. It also incurs an additional one billion ($10^{9}$) environment samples prior to policy training. All of our agents significantly outperform the extrinsic only baseline on the sparse reward tasks, demonstrating they are able to explore effectively while synthesizing their intrinsic rewards from online data alone.
@@
ELLM-BoW performs quite well on these tasks, closely tracking or matching Motif and ONI methods. However, in Section 5.2 we highlight a fundamental limitation of ELLM-BoW, namely its inability to capture complex semantic meaning, whereas ONI is capable thanks to its use of an LLM. It is worth noting that directly using ELLM-BoW as in Du et al. (2023a) without our episodic bonus completely fails, see Section C.2.
Game Progress of Intrinsic-Reward-Only Agents Results for all methods trained in the reward-free setting are shown in Figure 3(b). All of our ONI methods are able to make meaningful progress across all metrics. In particular, ONI-classification outperforms Motif on 3 out of 4 metrics. We include the results of Motif for reference, yet emphasize it is actually inapplicable in a truly reward-free setup since it assumes access to dense rewards.

### v4

@@ 1 Introduction
Reward functions are central to reinforcement learning (RL), and are often assumed to be given as part of the problem definition (Sutton & Barto, 2018). These functions are written to describe the task at hand, and often involve tradeoffs between ease of task definition and ease of policy optimization. For example, assigning a reward of +1+1 for solving the task and 0 otherwise is simple to define and accurately reflects the task goal, but is difficult to optimize due to providing zero gradients almost everywhere.
@@
In this work, we present ONI, a distributed online intrinsic reward and agent learning system — Figure˜1 overviews the high-level components. ONI removes the dependency on external datasets beyond the agent’s own experience and enables large-scale RL training with ease. ONI assumes access to captions of observations, similar to previous work Klissarov et al. (2023); Chu et al. (2023). The captions of collected observations are annotated online by an asynchronous LLM server, and both the policy and intrinsic reward model are simultaneously updated using the LLM’s feedback. Such a learning framework allows us to easily instantiate different algorithmic choices for synthesizing LLM feedback. Specifically, we explore three methods: the first one is retrieval-based and simply hashes the annotations; the second builds a binary classification model to distill the sentiment labels returned by the LLM; and the third sends pairs of captions to the LLM server for preference labeling and learns a ranking model, similar to Motif. We demonstrate that ONI is able to match Motif’s performance across a range of challenging, sparse rewards from the NetHack Learning Environment (NLE) (Küttler et al., 2020), solely using the agent’s gathered experience in a single, unified process. We also open-source our code to facilitate future work on LLM-based intrinsic motivation in settings where large offline datasets are not available.
@@ 2 Background
A key research question is how to define or learn the intrinsic reward $r^{\mathrm{int}}$. A first option is to manually define $r^{\mathrm{int}}$ based on task-specific goals, for example a measure of the distance between the agent’s current state and the goal state. However, handcrafting the intrinsic reward function can require significant domain knowledge and must be redone for each new task. A second option is to define $r^{\mathrm{int}}$ to measure some notion of observation novelty, which encourages the agent to systematically explore the environment. This can work well in smaller environments, but fails in ones that cannot be exhaustively explored in a tractable amount of time. A third class of methods, which we focus on in this work, leverage LLMs to automatically synthesize $r^{\mathrm{int}}$ to reflect prior knowledge about the task. We discuss all three classes of methods in more detail in Section˜4.
@@ 3.1 System design: distributed PPO with LLM annotations
Our core system, illustrated in Figure˜2, is built on top of the Sample Factory library v1.0 (Petrenko et al., 2020) and their asynchronous variant of proximal policy optimization (Schulman et al., 2017), referred to as APPO. APPO operates on a single machine and concurrently runs many environment instances while asynchronously updating policy and value estimates with the usual PPO rules, and adequately handles policy staleness and data transfers between these components. Concretely for NetHack, by running 480 environment instances APPO collects approximately 32k environment interactions per second on a Tesla A100-80GB GPU with 48 CPUs.
@@
For online reward learning via LLM annotations in this system [^1] The prior work by Klissarov et al. (2023) in the offline setting was able to connect their intrinsic reward function into Sample Factory’s APPO implementation with a few lines of code, as it just needed to load the PyTorch module of the learned reward function., we added 1) an LLM server hosted on a separate node, 2) an asynchronously running process that passes observation captions to the LLM server via HTTP request, 3) a hash table that stores the captions and corresponding LLM annotations, and 4) learning code that dynamically updates a reward model that learns on these annotations. Without being asynchronous, these components have the potential to block the concurrent execution and can reduce the throughput of the overall system, because calling into an LLM and updating the reward model are time-consuming. We added them in a way that retains most of the throughput (!), approximately 80-95% of the original: the average throughput is 30k environment interactions per second if we do not train any additional reward model to distill the LLM annotations, and 26k if a classification-based reward model is learned at the same time (see Section˜3.2 for the description of those approaches). The throughput of LLM annotation depends on the actual LLM and prompt we are using. For the reader’s reference, when hosting a LLaMA-3.1-8B server on a Tesla A100-80GB GPU using the prompt displayed in Appendix˜B, ONI annotates approximately 810k NetHack captions over the whole training process (2B environment steps, 19-20 hours). The rest of this section overviews the relevant components and how we have modified them.
@@
Our New Asynchronous LLM Worker and Remote LLM Server This worker has input and output queues that communicate back with the learner process, which we use in a way that does not block the main execution of the system. The LLM worker awaits new observations to label from the input queue, formats them into prompts, calls into an LLM, and returns the annotations in the output queue. Additionally, the prompt templates and other LLM options are configured by the LLM worker. We use an annotation and message format that support all the intrinsic reward labels that we consider in Section˜3.2. As Sample Factory already utilizes most of the free CPU and GPU capacity of the system, we opted to call into the LLM via an HTTP/REST interface rather than loading it in this process directly. The communication here adds minimal overhead, and also avoids the need to coordinate the shared memory of multiple GPUs between the main APPO code and LLM. Alongside every APPO run, we allocate an unloaded machine with a new instance of vLLM (Kwon et al., 2023b) to exclusively serve the queries from the RL run. Our system also allows using the same node for both processes if computational resources allow, which further reduces communication costs.
@@
To prevent overfitting the limited amount of annotations in the early stage of training, the training thread only continuously updates the reward model after we receive 2525k annotations. Before that, the feedback processing thread runs a few updates of the reward model every time we receive new annotations. Now that we have a flexible way of annotating messages and fitting a reward model on top of them, we turn to defining the types of intrinsic reward functions.
@@ 3.2 Intrinsic Reward Functions
If $c_{t}$ is unlabeled, it is placed into a last-in-first-out (LIFO) queue $Q$ managed by the LLM annotation process, and then sent to the LLM server. The LLM continuously processes elements $c_{i}$ from $Q$ and returns their labels $y_{i}$ to the data processing thread of the learner worker, where the pairs $(c_{i},y_{i})$ are added into $\mathcal{H}$. As described in Section˜3.1, the training thread and the feedback processing thread run asynchronously. Therefore, editing the hash table does not slow down policy training. This retrieval-based approach does not generalize to observations with unlabeled captions. However, the resulting intrinsic reward is simple and hyperparameter-free, and may work well when the set of captions belongs to a relatively small set.
@@
where $\mathbb{I}$ is the indicator function. We study the impact of $\eta$ in Section˜5.3. Unlike the previous approach, this method has the potential to generalize to observations whose captions are similar, but not identical, to the captions labeled by the LLM. However, like the previous approach, it will assign a same reward to observations which are slightly positive (such as finding a few gold pieces in NetHack) and very positive (finding hundreds of gold pieces or a rare artifact).
@@
where we use average log-likelihood when $y=\varnothing$ [^3] In the equivalent cross entropy minimization formulation, this amounts to using a uniform target $\frac{1}{2},\frac{1}{2}]$ when $y=\varnothing$. and use the Bradley-Terry model Bradley & Terry (1952) $P_{\phi}(o_{1}\succ o_{2})=\exp\big(r^{\mathrm{int}}_{\phi}(o_{1})\big)/\big[\exp\big({r^{\mathrm{int}}_{\phi}(o_{1})}\big)+\exp\big(r^{\mathrm{int}}_{\phi}(o_{2})\big)\big]$. Motif computes the mean $\mu_{\mathcal{D}}$, standard deviation $\sigma_{\mathcal{D}}$, and a fixed quantile $\nu_{\mathcal{D}}$ of $r^{\mathrm{int}}_{\phi}$ over the offline dataset of annotations $\mathcal{D}$. During RL training, it normalizes and thresholds $r^{\mathrm{int}}_{\phi}$ to give the reward
@@ 4 Related Work
Goal-conditioned Reward Design A different approach to reward function design is to define rewards as the distance between the agent’s current state and the goal. For example, one line of work learns a state embedding in a self-supervised manner which converts geodesic distances in the original space to Euclidean distances in feature space (Wu et al., 2019; Wang et al., 2021; Gomez et al., 2024). Another line of work of (Fan et al., 2022; Rocamonde et al., 2023; Adeniji et al., 2023; Kim et al., 2024) leverages pretrained image and text encoders, and defines rewards to be some measure of similarity between embeddings of visual observations and embeddings of textual task descriptions. Using a question generation and answering system, Carta et al. (2022) extract auxiliary objectives from the goal description and construct intrinsic rewards. An interesting combination of goal-conditioned and LLM-aided reward design is the ELLM approach introduced in Du et al. (2023b), which generates candidate goals and uses the distance in LLM embedding space to define the reward. This approach shares the limitations of the works of Kwon et al. (2023a); Chu et al. (2023) discussed in the previous section, in that it requires an LLM call for each agent observation, which becomes computationally infeasible in high-throughput settings involving billions of observations. We discuss more works that utilize LLM for RL broadly in Appendix˜D.
@@ 5 Experiments
Methods and Hyperparameters We instantiate ONI with the three reward functions described in Section˜3.2. We name the three approaches ONI-retrieval, ONI-classification and ONI-ranking, respectively. For policy learning, we use the Chaotic Dwarven GPT5 (CDGPT5) architecture (Myffili, 2021) used in prior work (Piterbarg et al., 2023; Kurenkov et al., 2023; Klissarov et al., 2023). Architecture details can be found in Section˜A.1. All the methods are trained with two billion $(2\times 10^{9})$ environment steps. We train ONI-classification with with classification threshold $\eta=0.5$, where the intrinsic reward coefficient is $\beta=0.1$ for the Score task and $\beta=0.4$ for all the other sparse-reward tasks and the reward-free agent. Similarly, ONI-retrieval uses $\beta=0.1$ for Score and $\beta=0.5$ for the others. ONI-ranking is trained with $\beta=0.05$ and $\nu_{\mathcal{N}}=0$. Full details of the training process can be found in Section˜A.2.
LLMs We use the LLaMA-3 herd of models (Dubey et al., 2024) as our LLMs. All prompts are listed in Appendix˜B. We initially reran the Motif baseline using the official code [^5] https://github.com/facebookresearch/motif and compared the performance of LLaMA-3.1-70B-Instruct and LLaMA-3.1-8B-Instruct (see Section˜C.1). We did not observe a significant difference in their performance on the Oracle or Score tasks. Therefore, we use LLaMA-3.1-8B-Instruct in our subsequent experiments to reduce computation.
Baselines We compare to 3 baselines: i) agents trained with extrinsic reward alone, ii) Motif (Klissarov et al., 2023), and iii) a variant of the ELLM algorithm Du et al. (2023a). It is not feasible to directly apply ELLM to our setting, since it requires an LLM call for each observation, where the total number of calls scales to the billions in our case. Therefore, we designed a more scalable variant which i) replaces the LLM embedding with a lightweight bag-of-words embedding  [^6] We also tried using Sentence Transformers (SBERT) for generating sentence embeddings, which led to an approximately 20x throughput drop: 28k vs 1.4k FPS on a Tesla V100-32GB GPU, when the sentence embeddings can be cached. If we recompute the sentence embedding in the forward pass every time, the throughput drop is approximately 100x. using FastText (Bojanowski et al., 2016), and ii) includes an episodic term in the intrinsic reward, as with Motif and ONI (see Section˜A.2). We call it ELLM-BoW and include more details in Appendix A.3. We note that Klissarov et al. (2023) found that other exploration methods based on novelty bonuses such as RND (Burda et al., 2019), NovelD (Zhang et al., 2021) and E3B (Henaff et al., 2022a) did not improve over the extrinsic reward baseline on these tasks, hence we do not include them.
@@ 5.1 Main Results
Task Performance We report the average performance and $95\%$ confidence intervals computed via standard errors over $5$ seeds in Figure˜3(a). The extrinsic reward agent performs reasonably well on the dense Score task, but completely fails on the others due to reward sparsity. We find that, despite not requiring any external data, our ONI-classification agent is able to match Motif on all tasks except Oracle, where it still performs well. We note that Motif requires pre-collecting data from the Score task even for sparse reward tasks—this assumes access to an additional dense reward function, which is an often unrealistic assumption. It also incurs an additional one billion ($10^{9}$) environment samples prior to policy training. All of our agents significantly outperform the extrinsic only baseline on the sparse reward tasks, demonstrating they are able to explore effectively while synthesizing their intrinsic rewards from online data alone.
@@
ELLM-BoW performs quite well on these tasks, closely tracking or matching Motif and ONI methods. However, in Section˜5.2 we highlight a fundamental limitation of ELLM-BoW, namely its inability to capture complex semantic meaning, whereas ONI is capable thanks to its use of an LLM. It is worth noting that directly using ELLM-BoW as in Du et al. (2023a) without our episodic bonus completely fails, see Section˜C.2.
Game Progress of Intrinsic-Reward-Only Agents Results for all methods trained in the reward-free setting are shown in Figure˜3(b). All of our ONI methods are able to make meaningful progress across all metrics. In particular, ONI-classification outperforms Motif on 3 out of 4 metrics. We include the results of Motif for reference, yet emphasize it is actually inapplicable in a truly reward-free setup since it assumes access to dense rewards.